what are the solutions of 9x^2+16=0

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

what are the solutions of 9x^2+16=0
-----
Factor using imaginary numbers::
(3x-4i)(3x+4i) = 0
---------------
x = (4/3)i or x = (-4/3)i
-------------
Cheers,
Stan H.
----------------

Answer by addingup(3677)

9x^2+16=0 = 9x^2= -16; √9x^2=√-16; 9x= 4i; x= 4i/9
Note: keep in mind that any time there is a negative under the radical sign, it comes out from underneath as the imaginary number "i" because the problem will contain a square root of -1, since i= √-1